APPROXIMATE THE VOLUME OF A TIN CAN 8 INCHES TALL AND WITH A DIAMETER OF 7 INCHES. USE 3.14 FOR PI. PROVIDE AN EXPLANATION WITH YOUR ANSWER
Answer by dabanfield(803)

The can is a right-circular cylinder so the formula for the volune is:
V = (pi)*(r^2)*h where r is the radius and h is the height of the can. Using 3.14 for pi, 8 for the height and 7/2=3.5 (diameter divided by 2) for the radius you have:
V = (3.14)*((3.5)^2)*8
You can now calculate the answer.